# Exploiting Observation Bias to Improve Matrix Completion

### v2

@@ Results.
As we can see from Fig. 4, the estimates from modified USVT are extremely biased when the algorithm is used with MNAR data. The estimates from MNN, however, appear to be minimally biased. Moreover, from Fig. 3, we can see that the estimates made by modified USVT are very sensitive to outliers in the data, while the estimates from MNN are not. We also observe from Table 2 that the $R^{2}$, MSE, MAE, and max error of MNN are far better than that of modified USVT, implying that MNN works significantly better on MNAR data in practice.
Figure 3: Distributions of test set estimates from MNN and modified USVT for real-world data.
Figure 4: Predictions vs true outcomes for real-world data using MNN (left) and modified USVT (right).
Algorithm $R^{2}$ MSE MAE Max Error
USVT $-26.0\pm.435$ $32.4\pm.392$ $4.95\pm.032$ $11.6\pm.053$
Table 2: $R^{2}$, MSE, MAE, and max error for matrix completion methods on real data.
Appendix A Distance Estimation
@@
Lemma A.1 (Distance estimation).
@@
$\hat{d}^{\mathbf{u}}_{ij}=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widehat{\widetilde{\mathbf{P}}}_{j,:}\right\rVert_{2}\qquad\text{and}\qquad d^{\mathbf{u}}_{ij}=\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}.$ (26)
@@
$\tilde{d}_{ij}^{\mathbf{u}}=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widetilde{\mathbf{P}}_{i,:}-\widetilde{\mathbf{P}}_{j,:}\right\rVert_{2}$ $=\sqrt{\frac{d}{n}}~{}\left\lVert\mathbf{V}(\mathbf{u}_{i}-\mathbf{u}_{j})\right\rVert_{2}.$ (29)
@@
[…]
@@
$\left|\tilde{d}_{ij}^{\mathbf{u}}-d^{\mathbf{u}}_{ij}\right|=\left|\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widetilde{\mathbf{P}}_{i,:}-\widetilde{\mathbf{P}}_{j,:}\right\rVert_{2}-\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}\right|\leq\max_{k\in[d]}\left|\sqrt{\frac{d}{n}}~{}\sigma_{k}\left(\mathbf{V}\right)-1\right|\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}$ (31)
where $\sigma_{\min}(\mathbf{X})$ (resp. $\sigma_{\max}(\mathbf{X})$) refers to the smallest (resp. largest) singular value of $\mathbf{X}$. Since $\sqrt{d}\cdot\mathbf{V}$ is a random matrix with i.i.d. isotropic rows, using Lemma H.5 and picking $t=\sqrt{\log(2/\delta)/c}$, we obtain that
@@
Therefore controlling the term $|\hat{d}^{\mathbf{u}}_{ij}-\tilde{d}^{\mathbf{u}}_{ij}|$ amounts to showing that $\|\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widetilde{\mathbf{P}}_{i,:}\|_{2}$ is small for all $i\in[n]$. To that end, we use Lemma H.2 to write
$\left\|\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widetilde{\mathbf{P}}_{i,:}\right\|_{2}^{2}\leq 2\cdot\left(\frac{\|\mathbf{E}\|_{\textup{op}}^{2}}{\sigma_{d}^{2}(\widetilde{\mathbf{P}})}\left(\|\mathbf{E}_{j,:}\|_{2}^{2}+\|\widetilde{\mathbf{P}}_{j,:}\|_{2}^{2}\right)+\|\mathbf{R}_{d}\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}\|_{2}^{2}\right),$ (36)
where we denote $\mathbf{E}=\widetilde{\mathbf{A}}-\widetilde{\mathbf{P}}=\mathbf{A}-\mathbf{P}$, and $\mathbf{Q}_{d}\mathbf{\Sigma}_{d}\mathbf{R}_{d}^{\top}$ the reduced SVD of $\widetilde{\mathbf{P}}$. Now, we will produce several bounds on the norms of the relevant matrices appearing in (36) separately.
(2.1) Upper bounding $\left\lVert\mathbf{E}\right\rVert_{\operatorname{op}}$. The entries of the noise matrix $\mathbf{E}$ are independent centered Bernoulli random variables, thus sub-gaussian. Using Lemma H.3 and choosing $t=\sqrt{\log(2/\delta)/c}$ again, we obtain that the following
@@
We will bound this quantity using Lemma H.7. Towards showing that $\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}$ is a sub-gaussian random vector, consider any fixed $\mathbf{v}\in\mathcal{S}^{d-1}$. Then, denoting $\mathbf{R}_{d}=[r_{ki}]_{k\in[n],i\in[d]}$,
@@
Therefore, $\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}$ is a $(1/4)$-sub-gaussian random vector. Using Lemma H.7 with $t=\sqrt{\frac{d}{2}\log\left\lparen\frac{2n}{\delta}\right\rparen}$ and union bounding over all rows of $\mathbf{E}$ yields
@@
(2.6) Putting everything together. Starting from (36), and using the bounds obtained in (37), (40), (42), (43), and (53), we have by substitution that
@@
$\leq C\cdot\left\lparen\frac{C^{\prime}d^{2}}{\rho_{n}^{2}}+\frac{d}{2}\log\left\lparen\frac{2n}{\delta}\right\rparen\right\rparen$ ) ) (56)
$\leq\frac{Cd^{2}}{\rho_{n}^{2}}\log\left\lparen\frac{2n}{\delta}\right\rparen.$ ) . (57)
Scaling appropriately according to (26), and using the union bound, then yields
@@
provided $n\geq cd\log(en/\delta)$, for some universal constants $c,C>0$. ∎
Appendix B Clustering
Coverage. A key phenomenon that underpins our results is dense coverage of the latent feature space by users and items. Towards guaranteeing this, let $\Sigma_{n}\subset\mathcal{S}^{d-1}$ be a minimal $\varepsilon_{n}$-net of $\mathcal{S}^{d-1}$, with the properties that $\left\lVert\mathbf{z}-\mathbf{z}^{\prime}\right\rVert_{2}>\varepsilon_{n}$ for distinct $\mathbf{z},\mathbf{z}^{\prime}\in\Sigma_{n}$ and for any $\mathbf{x}\in\mathcal{S}^{d-1}$ there exists $\mathbf{z}\in\Sigma_{n}$ such that $\left\lVert\mathbf{x}-\mathbf{z}\right\rVert_{2}\leq\varepsilon_{n}$. It is widely known that $|\Sigma_{n}|\leq(1+2/\varepsilon_{n})^{d}\leq(3/\varepsilon_{n})^{d}$ where the latter inequality assumes $\varepsilon_{n}\leq 1$, which we enforce.
Lemma B.1 (Coverage).
@@ Appendix B Clustering
$\mathbb{P}\left(\min_{\mathbf{z}\in\Sigma_{n}}\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\geq\frac{n}{8}\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}\right)\geq 1-\delta,$ (60)
@@
$\mathbb{P}\left(\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\leq\frac{n}{2}\,\mathbb{P}(\mathbf{u}_{1}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}))\right)\leq\exp\left(-\frac{n}{8}\,\mathbb{P}(\mathbf{u}_{1}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}))\right)$ (62)
Now, by using Lemma H.1 we have
@@
$\sin\Phi=\varepsilon_{n}\sqrt{1-\varepsilon_{n}^{2}/4}\geq\varepsilon_{n}/2.$ (64)
Thus we can write
$\mathbb{P}(\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}))\geq\frac{1}{4}\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}.$ (65)
@@
$\mathbb{P}\left(\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\leq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right)\leq\exp\left(-\frac{n}{32}\,\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}\right).$ ) . (66)
@@
$\mathbb{P}\left(\min_{\mathbf{z}\in\Sigma_{n}}\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\leq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right)$ $\leq\left(\frac{3}{\varepsilon_{n}}\right)^{d}\exp\left(-\frac{n}{32}\,\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}\right)$ ) (67)
$\leq\delta,$ (68)
@@
Observations. Given that the latent feature space is densely covered by both users and items, we now argue that all pairs of user and item clusters – as defined by $\Sigma_{n}$, rather than the central users and items – that are reasonably close will generate a guaranteed number of observations.
Lemma B.2 (Observations).
Consider $\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}$ for which $\left\lVert\mathbf{z}_{i}-\mathbf{z}_{j}\right\rVert_{2}\leq\sqrt{2}+c\varepsilon_{n}$ for some absolute constant $c>0$, where we assume that $\varepsilon_{n}<(\sqrt{3}-\sqrt{2})/(2+c)$, and let the set of observations between these user and item clusters be denoted by
$\mathcal{Y}_{\varepsilon_{n}}(\mathbf{z}_{i},\mathbf{z}_{j})=\{y_{kl}:a_{kl}=1,\mathbf{u}_{k}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i}),\mathbf{v}_{l}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{j})\}.$ (70)
Then, for all $\delta\in(0,1)$,
$\mathbb{P}\left(\min_{\begin{subarray}{c}\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}:\\ \|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+c\varepsilon_{n}\end{subarray}}\left\lvert\mathcal{Y}_{\varepsilon_{n}}(\mathbf{z}_{i},\mathbf{z}_{j})\right\rvert\geq cn^{2}\rho_{n}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2}\right)\geq 1-\delta,$ (71)
as long as $\varepsilon_{n}$ satisfies

### v3

@@ Results.
The results (see Table 3) suggest that the performance of MNN is comparable to that of SNN for relatively small values of $n$, namely $n=100$ and $n=200$, with SNN slightly outperforming MNN on the MAE metric. However, we observed that as $n$ grows large the computational performance of SNN degrades rapidly in comparison with our implementation of MNN. This is to be expected because SNN needs to solve a biclique search problem which is known to be a computationally hard problem.
Comparison with SNN on Real-data. We tried to also compare with SNN on the Glance dataset. In this dataset, the number of users is $1305$ and the number of content items is $1471$. Unfortunately, SNN was not able to finish running in a reasonable time. Indeed, even on the synthetically generated dataset with only $300$ users and $300$ items, SNN could not finish running within 24 hours as reported in the table because it is not computationally efficient (see section 4.2. in [4]). To be able to compare the with our algorithm, we would require a distributed implementation of SNN, which is beyond the scope of this work. This also means that our proposed algorithm is meaningful.
Table 3: MSE, MAE, and runtime for MNN and SNN on synthetic datasets (average $\pm$ standard deviation across 10 experimental repeats). --- - means the method did not complete within $24$ hours.
Algorithm $n=100$ $n=200$ $n=300$ $n=1000$
RMSE MNN $8.96\pm 0.18$ $8.93\pm 0.06$ $\mathbf{8.44\pm 0.12}$ $\mathbf{6.16\pm 0.06}$
SNN $\mathbf{7.60\pm 0.17}$ $\mathbf{7.98\pm 0.07}$
MAE MNN $7.10\pm 0.23$ $7.06\pm 0.05$ $\mathbf{6.63\pm 0.14}$ $\mathbf{4.38\pm 0.05}$
SNN $\mathbf{4.53\pm 0.08}$ $\mathbf{4.78\pm 0.03}$
Run Time MNN $\leq$ 10 s $\leq$ 10 s $\leq$ 10 s $\leq$ 20 s
SNN 1 min 43 s 40 min 19 s $\geq$ 24 h $\geq$ 24 h
6.4 Practical implementation of MNN
As highlighted before, in the experiments, we utilize a practical implementation of the proposed MNN algorithm where (i) $\mathbf{A}$ is used in place of $\tilde{\mathbf{A}}$ in the distance estimation step and (ii) the matrix completion from the clustered outcome matrix is done using SVD with known rank. Using such a practical implementation of the algorithm acts a proof of concept – letting us determine the added value of the extra information gained from the two stage approach.
Regarding clustering in (i).
Typically one may perform clustering by either specifying the number of clusters or specifying a threshold distance as described in the proposed MNN in Section 3. We use the subroutine the routine AgglomerativeClustering from scikit.cluster which allows us to perform clustering in either ways. However, we still need to specify the number of clusters or the threshold distance that both depend on $\varepsilon_{n}$. In view of our discussion on the coverage property (see Appendix A.2), a conservative choice on the number of clusters is $(3/\varepsilon_{n})^{d}$, while a conservative choice of a distance threshold is $6\varepsilon_{n})$ as per the description of MNN in Section LABEL:Alg:mnn. Note that according to Theorem 1, $\varepsilon_{n}=2n^{-(2-\beta)/(d-1)}$ and the value of $\beta$ may not be known, but we may simply estimate as was the case for the Glance dataset. Practically, we observed that both ways of clustering perform similarly, but we had to tune the constants in front of $\varepsilon_{n}$ appropriately. In particular, when we performed clustering based on the number of clusters, we selected the number of cluster within the values $\{\sqrt{n}/2,\sqrt{n},2\sqrt{n},5\sqrt{n}\}$, and when we performed clustering based on a distance threshold, we selected the thresholds within the values $\{\varepsilon_{n}/5,\varepsilon_{n}/6,\varepsilon_{n}/7,\dots\varepsilon_{n}/90\}$. The selection process was done via cross-validation.
Regarding imputation in (ii).
Note that the only part of the proposed MNN that may be practically inefficient is the imputation subroutine as it requires to search over the space of subsets of $[n]$ of size $d$, and such space is of size (${n\choose d}$. Thus, the imputation subroutine may have a computational complexity that is of order $n^{d\log(d)}$ which is still polynomial in $n$ but may not be practical for large values of $n$. Devising a practical efficient procedure to perform imputation under our model while enjoying entry-wise guarantees is an interesting open question that we leave for future work. For practical purposes, we recommend to use a standard matrix completion algorithm for this second phase such as USVT, ALS or nuclear norm penalization based estimators as we do. However, note that these methods are not proven to enjoy entry-wise guarantees under our setting which we find to be an interesting question that we leave for future work.
7 Conclusion
We proposed MNN a matrix completion method suitable for MNAR data models where there is shared information between the outcome and observation pattern. The proposed method provably exploits this shared information to improve predictions. The desirable theoretical properties and empirical performance of MNN highlight the value of modeling these shared latent factors. From a statistical learning perspective, we have shown that the parametric rate for supervised learning algorithms such as nearest-neighbor-style method can almost be achieved even when (i) covariates are unobserved, and (ii) observations are sampled in a biased manner; this could be of independent interest. Finally, we note that valuable future directions of research include (i) devising practical matrix completion methods for generic MNAR data models that are as statistically efficient as MNN, and (ii) potentially relaxing conditional independence assumption between different entries of the observation pattern.
Appendix A Proof of Proposition 2
A.1 Distance Estimation
@@
Lemma 2 (Distance estimation).
@@
$\hat{d}^{\mathbf{u}}_{ij}$ $=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widehat{\widetilde{\mathbf{P}}}_{j,:}\right\rVert_{2},$
$d^{\mathbf{u}}_{ij}$ $=\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}.$ (43)
@@
$\tilde{d}_{ij}^{\mathbf{u}}=\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widetilde{\mathbf{P}}_{i,:}-\widetilde{\mathbf{P}}_{j,:}\right\rVert_{2}$ $=\sqrt{\frac{d}{n}}~{}\left\lVert\mathbf{V}(\mathbf{u}_{i}-\mathbf{u}_{j})\right\rVert_{2}.$
@@
$\sqrt{\frac{d}{n}}\sigma_{\min}\left\lparen\mathbf{V}\right\rparen\left\lVert\mathbf{u}_{i}-\mathbf{u}_{j}\right\rVert_{2}\leq\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widetilde{\mathbf{P}}_{i,:}-\widetilde{\mathbf{P}}_{j,:}\right\rVert_{2}$
$\qquad\qquad\qquad\qquad\quad\leq\sqrt{\frac{d}{n}}\sigma_{\max}\left\lparen\mathbf{V}\right\rparen\left\lVert\mathbf{u}_{i}-\mathbf{u}_{j}\right\rVert_{2},$
@@
$\left|\tilde{d}_{ij}^{\mathbf{u}}-d^{\mathbf{u}}_{ij}\right|$ $=\left|\frac{2}{\rho_{n}}\sqrt{\frac{d}{n}}~{}\left\lVert\widetilde{\mathbf{P}}_{i,:}-\widetilde{\mathbf{P}}_{j,:}\right\rVert_{2}-\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}\right|$
$\leq\max_{k\in[d]}\left|\sqrt{\frac{d}{n}}~{}\sigma_{k}\left(\mathbf{V}\right)-1\right|\|\mathbf{u}_{i}-\mathbf{u}_{j}\|_{2}$
where $\sigma_{\min}(\mathbf{X})$ (resp. $\sigma_{\max}(\mathbf{X})$) refers to the smallest (resp. largest) singular value of $\mathbf{X}$. Since $\sqrt{d}\cdot\mathbf{V}$ is a random matrix with i.i.d. isotropic rows, using Lemma 14 and picking $t=\sqrt{\log(2/\delta)/c}$, we obtain that
@@
Therefore controlling the term $|\hat{d}^{\mathbf{u}}_{ij}-\tilde{d}^{\mathbf{u}}_{ij}|$ amounts to showing that $\|\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widetilde{\mathbf{P}}_{i,:}\|_{2}$ is small for all $i\in[n]$. To that end, we use Lemma 11 to write
$\left\|\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widetilde{\mathbf{P}}_{i,:}\right\|_{2}^{2}\leq$
$\quad 2\left(\frac{\|\mathbf{E}\|_{\textup{op}}^{2}}{\sigma_{d}^{2}(\widetilde{\mathbf{P}})}\left(\|\mathbf{E}_{j,:}\|_{2}^{2}+\|\widetilde{\mathbf{P}}_{j,:}\|_{2}^{2}\right)+\|\mathbf{R}_{d}\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}\|_{2}^{2}\right),$ (50)
where we denote $\mathbf{E}=\widetilde{\mathbf{A}}-\widetilde{\mathbf{P}}=\mathbf{A}-\mathbf{P}$, and $\mathbf{Q}_{d}\mathbf{\Sigma}_{d}\mathbf{R}_{d}^{\top}$ the reduced SVD of $\widetilde{\mathbf{P}}$. Now, we will produce several bounds on the norms of the relevant matrices appearing in (A.1) separately.
(2.1) Upper bounding $\left\lVert\mathbf{E}\right\rVert_{\operatorname{op}}$. The entries of the noise matrix $\mathbf{E}$ are independent centered Bernoulli random variables, thus sub-gaussian. Using Lemma 12 and choosing $t=\sqrt{\log(2/\delta)/c}$ again, we obtain that the following
@@
We will bound this quantity using Lemma 16. Towards showing that $\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}$ is a sub-gaussian random vector, consider any fixed $\mathbf{v}\in\mathcal{S}^{d-1}$. Then, denoting $\mathbf{R}_{d}=[r_{ki}]_{k\in[n],i\in[d]}$,
@@
Therefore, $\mathbf{R}_{d}^{\top}\mathbf{E}_{j,:}$ is a $(1/4)$-sub-gaussian random vector. Using Lemma 16 with $t=\sqrt{\frac{d}{2}\log\left\lparen\frac{2n}{\delta}\right\rparen}$ and union bounding over all rows of $\mathbf{E}$ yields
@@
(2.6) Putting everything together. Starting from (A.1), and using the bounds obtained in (51), (54), (56), (57), and (67), we have by substitution that
$\left\lVert\widehat{\widetilde{\mathbf{P}}}_{i,:}-\widetilde{\mathbf{P}}_{i,:}\right\rVert_{2}^{2}$
@@
$\quad\leq C\cdot\left\lparen\frac{C^{\prime}d^{2}}{\rho_{n}^{2}}+\frac{d}{2}\log\left\lparen\frac{2n}{\delta}\right\rparen\right\rparen$ ) )
$\quad\leq\frac{Cd^{2}}{\rho_{n}^{2}}\log\left\lparen\frac{2n}{\delta}\right\rparen.$ ) .
Scaling appropriately according to (A.1), and using the union bound, then yields
@@
provided $n\geq cd\log(en/\delta)$, for some universal constants $c,C>0$.
A.2 Clustering
A.2.1 Coverage.
A key phenomenon that underpins our results is dense coverage of the latent feature space by users and items. Towards guaranteeing this, let $\Sigma_{n}\subset\mathcal{S}^{d-1}$ be a minimal $\varepsilon_{n}$-net of $\mathcal{S}^{d-1}$, with the properties that $\left\lVert\mathbf{z}-\mathbf{z}^{\prime}\right\rVert_{2}>\varepsilon_{n}$ for distinct $\mathbf{z},\mathbf{z}^{\prime}\in\Sigma_{n}$ and for any $\mathbf{x}\in\mathcal{S}^{d-1}$ there exists $\mathbf{z}\in\Sigma_{n}$ such that $\left\lVert\mathbf{x}-\mathbf{z}\right\rVert_{2}\leq\varepsilon_{n}$. It is widely known that $|\Sigma_{n}|\leq(1+2/\varepsilon_{n})^{d}\leq(3/\varepsilon_{n})^{d}$ where the latter inequality assumes $\varepsilon_{n}\leq 1$, which we enforce.
Lemma 3 (Coverage).
@@ A.2 Clustering
$\mathbb{P}\Big{(}\min_{\mathbf{z}\in\Sigma_{n}}\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\geq\frac{n}{8}{\Big{(}\frac{\varepsilon_{n}}{2}\Big{)}}^{d-1}\Big{)}$ $\geq 1-\delta,$ (70)
@@
$\mathbb{P}\Big{(}\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\leq\frac{n}{2}\,\mathbb{P}(\mathbf{u}_{1}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}))\Big{)}$
$\qquad\leq\exp\Big{(}-\frac{n}{8}\,\mathbb{P}(\mathbf{u}_{1}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}))\Big{)}.$ (72)
Now, by using Lemma 10 we have
@@
$\sin\Phi=\varepsilon_{n}\sqrt{1-\varepsilon_{n}^{2}/4}\geq\varepsilon_{n}/2.$
Thus, we can write
$\mathbb{P}(\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}))\geq\frac{1}{4}\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}.$
@@
$\mathbb{P}\left(\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\leq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right)$
$\qquad\qquad\qquad\quad\quad\leq\exp\left(-\frac{n}{32}\,\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}\right).$ ) .
@@
$\mathbb{P}\left(\min_{\mathbf{z}\in\Sigma_{n}}\sum_{i=1}^{n}\mathds{1}\{\mathbf{u}_{i}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z})\}\leq\frac{n}{8}\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\right)$
$\qquad\qquad\qquad\quad\quad\leq\left(\frac{3}{\varepsilon_{n}}\right)^{d}\exp\left(-\frac{n}{32}\,\left\lparen\frac{\varepsilon_{n}}{2}\right\rparen^{d-1}\right)$ )
$\qquad\qquad\qquad\quad\quad\leq\delta,$
@@
A.2.2 Observations
Given that the latent feature space is densely covered by both users and items, we now argue that all pairs of user and item clusters – as defined by $\Sigma_{n}$, rather than the central users and items – that are reasonably close will generate a guaranteed number of observations.
Lemma 4 (Observations).
Consider $\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}$ for which $\left\lVert\mathbf{z}_{i}-\mathbf{z}_{j}\right\rVert_{2}\leq\sqrt{2}+c\varepsilon_{n}$ for some absolute constant $c>0$, where we assume that $\varepsilon_{n}<(\sqrt{3}-\sqrt{2})/(2+c)$, and let the set of observations between these user and item clusters be denoted by $\mathcal{Y}_{\varepsilon_{n}}^{(\mathbf{z}_{i},\mathbf{z}_{j})}=\{y_{kl}:a_{kl}=1,\mathbf{u}_{k}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{i}),\mathbf{v}_{l}\in\mathcal{B}_{\varepsilon_{n}}(\mathbf{z}_{j})\}$. Then, for all $\delta\in(0,1)$, the event
$\min_{\begin{subarray}{c}\mathbf{z}_{i},\mathbf{z}_{j}\in\Sigma_{n}:\\ \|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}\leq\sqrt{2}+c\varepsilon_{n}\end{subarray}}\left\lvert\mathcal{Y}_{\varepsilon_{n}}^{(\mathbf{z}_{i},\mathbf{z}_{j})}\right\rvert\geq cn^{2}\rho_{n}\left(\frac{\varepsilon_{n}}{2}\right)^{2d-2},$
holds with probability at least $1-\delta$, as long as $\varepsilon_{n}$ satisfies
$\left(\frac{\varepsilon_{n}}{2}\right)^{d-1}\geq\frac{C}{n}\left(\sqrt{\frac{1}{\rho_{n}}\log\left(\frac{e\,n}{\delta}\right)}+\log\left(\frac{e\,n}{\delta}\right)\right)$